# Estimating Privacy Leakage of Augmented Contextual Knowledge in Language Models

### v2

@@ Abstract
[…]
@@ 1 Introduction
Language Models (LMs) can rely on two sources of knowledge during generation: (1) parameteric knowledge, which is information from the LM’s pre-training corpora encoded within the model parameters Devlin et al. (2018); Radford et al. (2019); Petroni et al. (2019); (2) contextual knowledge, which is additional information passed into the input prompt Kwiatkowski et al. (2019); Joshi et al. (2017). For certain downstream tasks, such as question-answering, it is essential to augment prompts containing a question/instruction with relevant context for LMs. However, a recent concern is that both the parametric and contextual knowledge may contain private information. Prior work has shown that privacy leakage of parametric knowledge often occurs from memorized pre-training data Carlini et al. (2019). On the other hand, we focus on the privacy leakage of augmented contexts, which can occur when an LM regurgitates them Wang et al. (2023); Priyanshu et al. (2023).
@@
Consider the example shown in Figure 1. An augmented context contains John Doe’s address, and a user queries an LM asking for John Doe’s address. If the output of the LM contains John Doe’s address, then the straightforward approach of comparing the output against the augmented context would suggest there was privacy leakage from the context. This privacy evaluation was performed by prior works that studied data extraction attacks in RAG systems by prompting LMs to regurgitate the context Zeng et al. (2024); Qi et al. (2024). However, suppose we re-query the LM but remove (or mask out) John Doe’s address from the context. If the LM still outputs John Doe’s address, then surely the privacy leakage must derive from the LM’s parametric knowledge. Hence, assuming that augmented contexts are not contained in the LM’s parametric knowledge, which may not hold in practice Golchin and Surdeanu (2023); Deng et al. (2023); Jiang et al. (2024), can overestimate the privacy leakage. Indeed, our results in Section 4.2 demonstrates this. Thus, accurately measuring context privacy leakage requires consideration of the existing parametric knowledge of the LM.
Extensive research has examined how factors such as model size, prompt length, and training order contribute to the memorization and subsequent privacy leakage of an LM’s parametric knowledge Carlini et al. (2021, 2022); Biderman et al. (2024); Lesci et al. (2024). However, there is a lack of understanding regarding the factors that cause privacy leakage of contextual knowledge. This is challenging as it involves separating the contributions of an LLM’s parametric knowledge from the augmented context Longpre et al. (2021); Du et al. (2024), which has implications for solutions that adopt publicly pre-trained LMs to preserve privacy of contexts Utpala et al. (2023); Meisenbacher et al. (2024); Flemings and Annavaram (2024).
@@
  • We propose context influence, a metric to principally quantify the privacy leakage of contextual information by measuring the output difference with and without a subset of the context, exemplified in Table 1 with uni-grams. Context influence follows the analysis of differential privacy Dwork (2006), a widely-adopted privacy notion.
  • Then, using a slight reformulation of Context-aware Decoding Shi et al. (2023), we show that context privacy leakage can be affected (1) explicitly when amplifying/deamplifying contextual knowledge during decoding, and (2) implicitly when contextual knowledge is out-of-distribution with respect to parametric knowledge.
@@ 2.1 Privacy
To understand the privacy leakage of contextual information during decoding, we draw inspiration from Differential Privacy (DP) Dwork (2006); Dwork et al. (2014), a strong privacy notion that gives a provable guarantee on the information leakage. We state the definition below.
Definition 2.1 (Pure Differential Privacy (DP) Feldman and Zrnic (2021)).
@@
Definition 2.2 (Ex-post per-instance differential Privacy Redberg and Wang (2021)).
@@ 2.2 Context-aware Decoding
Satisfying the DP definition requires controlling the amount of privacy that can be leaked by the context during generation. To this end, we borrow ideas from prior work that focused on amplifying contextual information by utilizing Pointwise Mutual Information (PMI) to measure the LM’s dependence on the context, then applying this measurement to the decoding process to explicitly steer the LM’s focus on the context Van der Poel et al. (2022); Shi et al. (2023). The goal of these prior works does not involve privacy; rather, they are focused on reducing hallucinations by forcing the model to focus more on the context. However, these prior works provide the appropriate foundation to build on for estimating privacy from the context. PMI is defined as
@@
This formulation, known as Context-aware Decoding (CAD) Shi et al. (2023), helps the LM focus on the context.
@@ 3.1 Motivation
As we argued in the Introduction section, only comparing the LM’s output to the augmented context is insufficient for measuring context privacy leakage. Alternatively, one could estimate the privacy risk by performing Membership Inference Attacks (MIAs) Shokri et al. (2017); Jagielski et al. (2020). However, this requires instantiating an attacker, which could severely underestimate the privacy leakage, and MIAs have been shown to be oftentimes ineffective on LMs Duan et al. (2024).
Instead, we take a different approach by utilizing the privacy analysis from DP, which provides a strong guarantee that bounds the privacy leakage to any dataset $D$, individual $d_{i}$, and output events $E$. In particular, we follow the observation that sampling-based decoding naturally satisfies the randomized output requirement of differential privacy with respect to the context Flemings et al. (2024). Moreover, the neighboring definition of DP allows us to separate the contribution from parametric and contextual knowledge, by comparing the output probability distributions with and without subsets of the context. However, the $\epsilon$ bound from Eq. 1 does not provide much insights into the privacy risks of the augmented context, since $\epsilon$ is independent of the context. By using ex-post per-instance DP, we can directly calculate a privacy loss $\epsilon(o,D,D\setminus\{d_{i}\})$ that depends on important parameters– such as the context $D$, neighboring contexts $D\setminus\{d_{i}\}$, generated token $o$– which acts as a privacy auditing tool to analyze how these parameters affect the context privacy leakage.
@@ 3.2 Context Influence
$|\underbrace{\log p_{\theta}(y_{t}|D,\mathbf{x},\mathbf{y}_{<t})}_{\begin{subarray}{c}\text{output probability of $y_{t}$}\\ \text{given the context $D$}\end{subarray}}\!-\!\underbrace{\log p_{\theta}(y_{t}|D\!\setminus\!D_{i,n},\mathbf{x},\mathbf{y}_{<t})}_{\begin{subarray}{c}\text{output probability of $y_{t}$ with }\\ \text{$D_{i,n}$ removed from context}\end{subarray}}|$
The use of $i$-th token $n$-grams generalizes the privacy level [^1] We restrict the possible substrings to $n$-grams because they still produce natural text (contiguous substrings) while providing a granular measurement of the context influence., where $n=1$ roughly corresponds to word-level privacy Xu et al. (2020); Feyisetan et al. (2020) while $n=|D|$ corresponds to document-level privacy Mattern et al. (2022); Utpala et al. (2023) denoted as $D_{i,n}=D$. Rather than bounding the absolute value of the logs-odds ratio, as is done with DP, definition 3.1 directly uses this quantity as a lower-bound estimate of the privacy leakage of the $i$-th token $n$-gram $D_{i,n}$ for a fixed context $D$ when releasing the next token $y_{t}$. From an adversarial perspective, Definition 3.1 describes how confidently an attacker could infer whether the $i$-th token $n$-gram $D_{i,n}$ is part of the context $D$.
@@
To measure the total context influence over an entire generation $\mathbf{y}$, we simply sum the context influence for each generated token $y_{t}$ which is analogous to the basic composition property of DP Dwork et al. (2014).
@@
One last technicality: context influence only works for sampling-based algorithms, such as temperature sampling Ackley et al. (1985), and not for greedy decoding algorithms, such as argmax. However, top-p Holtzman et al. (2019) or top-k sampling Fan et al. (2018) can cause potential errors in the context influence calculation unless the selected indices are equal for both $D$ and $D\setminus D_{i,n}$. For the remainder of our work, we focus only on temperature sampling. Specifically, we generate the responses $\mathbf{y}$ to be used for measuring context influence by using a slight reformulation of CAD (Eq. 3), giving us granular control over how much of the contextual knowledge is used during decoding. We introduce Context Influence Decoding (CID) below:
@@
where $\sigma$ is the softmax function and $T$ is the temperature parameter where $T>1$ resulting in a more uniform distribution (i.e. higher entropy) and $10<T<$ 1 forcing a sharper output distribution.
@@ 3.3 Privacy Leakage with Context Influence
for a fixed $n$. The proof can be found in Appendix B. Meaning regardless of the context $D$, the previously generated tokens $\mathbf{y}_{<t}$, and the next token $y_{t}$, the $i$-th token $n$-gram $D_{i,n}$ influences the LM $\theta$ by at most $\epsilon$; hence, DP bounds the privacy leakage with a context-independent value. For our work, we want to explicitly measure how the privacy leakage changes with respect to the aforementioned variables, hence why chose the analysis of ex-post per-instance DP. This gives a guarantee that the privacy leakage when releasing $y_{t}$ is at least $\tau_{i,n}(\overline{p}_{\theta,\lambda},D,\mathbf{x},\mathbf{y}_{<t},y_{t})$, which follows the more practical direction of privacy auditing Jagielski et al. (2020).
@@
  • The context $D$ is out-of-distribution with respect to the LM’s parametric knowledge $\theta$, and the subset of contextual information we are interested in is sufficiently large (i.e. large $n$), both maximizing the difference of PMIs. Various factors– including the type of contextual $(D,\mathbf{x})$ and parametric knowledge $\theta$, model size $|\theta|$, and context size $|D|$– can cause the contextual and parametric knowledge to diverge, which we experimentally analyze in Section 4.3. Additionally, in Section 4.4, we compare how different subsets of contextual knowledge influence an LM.
@@
Model Decoding $\lambda$ $\hat{\tau}_{|D|}(\overline{p}_{\theta,\lambda})$ Repeat Prompts Rouge Prompts $\hat{\tau}_{|D|}(\overline{p}_{\theta,\lambda})$ Repeat Prompts Rouge Prompts
@@ 4.1 Experimental Setup
Datasets. We perform our experimental evaluations on two open-ended generation tasks: summarization via CNN-DM See et al. (2017), a collection of English news articles written by journalists at CNN and the Daily Mail, and long-form question-answering via PubMedQA Jin et al. (2019), a dataset from the biomedical domain and contexts available. Appendix C contains example prompts used for context influence. Each context document is truncated by $2048$ and $1024$ for PubMedQA and CNN-DM, respectively.
Metrics. We evaluate the generation quality along two dimensions: similarity and faithfulness. For similarity, we employed F1 ROUGE-L Lin (2004) and F1 BERTScore Zhang et al. (2019) to measure lexical and semantic similarity between the response and the reference, respectively. For faithfulness, we used FactKB Feng et al. (2023) to measure the faithfulness of the response to the context. Our calculation of context influence uses the empirical estimator in Eq. 5 with CID using sampled contexts from CNN-DM and PubMedQA.
Models. Since context influence requires access to the entire model output distribution, we use open-source models due to closed-source ones restricting the output logits. We used OPT 1.3B Zhang et al. (2022), GPT-Neo 1.3B Black et al. (2021), LLaMA 3 8B and LLaMA 3 8B IT (Instruct) Dubey et al. (2024), and Gemma 2 9B (Instruct) Team et al. (2024). We set temperature parameter $T=0.8$, the response length to at most 50 tokens, and the number of responses for each dataset is $N=1000$.
@@ 4.2 Context Influence on Input Regurgitation
First, we demonstrate how our context influence metrics offer improvements over directly comparing LLM output with augmented contexts. Following the untargeted attack evaluations by Zeng et al. (2024) we report:
@@
Next, we show that Repeat Prompts and Rouge Prompts can erroneously indicate privacy leakage from augmented PubMedQA abstracts if the LLM’s parametric knowledge already includes PubMed abstracts. For this, we compare OPT 1.3B and GPT-Neo 1.3B, models with the same number of parameters and similarly follow the GPT-3 architecture Brown et al. (2020). GPT-Neo was pre-trained on The Pile dataset Gao et al. (2020), which contains PubMed abstracts. In contrast, OPT was trained on a subset of The Pile that excludes PubMed abstracts Zhang et al. (2022). Therefore, we expect OPT 1.3B to show greater context privacy leakage, as it likely lacks PubMed abstracts in its parametric knowledge and must rely more on the augmented PubMedQA contexts.
@@ 4.3 Factors Contributing to Context Influence
Parametric knowledge $\theta$. We compare context influence on multiple different parametric knowledge sources. Our results indicate that the choice of parametric knowledge can have a substantial affect on the context influence. In particular, we saw from Section 4.2 how just the inclusion of PubMed data in the pre-training data can effectively decrease the context influence of PubMedQA abstracts, creating a false sense of security that GPT-Neo preserves the privacy of the PubMedQA abstracts better than OPT. Hence, this raises an important caveat that smaller context influence does not necessarily imply better context privacy, as one must consider the public data used for pre-training Tramèr et al. (2022).
@@
Context size $|D|$. Additionally, we measured the effect of the context size $|D|$ on context influence for CID using OPT-1.3B. In this setup, we restrict the model to only the first $|D|$ tokens of context for generation and calculating context influence. Shown in Figure 2(b), we observe that when the context is extremely small (32$\leq$ 32)≤ 32 ), then the LM is substantially less influenced by the context. The context may not contain enough relevant information to help the model, and hence, it must rely on its parametric knowledge. However, as we increase the context size from 32 to 256, the model becomes more influenced by the context. After $|D|\geq 256$, the model maintains a relatively constant level of context influence. Hence, truncating the context has marginal affect on context privacy unless the size of the context is substantially reduced.

### v3

@@ Abstract
Language models (LMs) rely on their parametric knowledge augmented with relevant contextual knowledge for certain tasks, such as question answering. However, the contextual knowledge can contain private information that may be leaked when answering queries, and estimating this privacy leakage is not well understood. A straightforward approach of directly comparing an LM’s output to the contexts can overestimate the privacy risk, since the LM’s parametric knowledge might already contain the augmented contextual knowledge. To this end, we introduce context influence, a metric that builds on differential privacy, a widely-adopted privacy notion, to estimate the privacy leakage of contextual knowledge during decoding. Our approach effectively measures how each subset of the context influences an LM’s response while separating the specific parametric knowledge of the LM. Using our context influence metric, we demonstrate that context privacy leakage occurs when contextual knowledge is out of distribution with respect to parametric knowledge. Moreover, we experimentally demonstrate how context influence properly attributes the privacy leakage to augmented contexts, and we evaluate how factors– such as model size, context size, generation position, etc.– affect context privacy leakage. The practical implications of our results will inform practitioners of the privacy risk associated with augmented contextual knowledge. https://github.com/james-flemings/context_influence
@@ 1 Introduction
Language Models (LMs) can rely on two sources of knowledge during generation: (1) parameteric knowledge, which is information from the LM’s pre-training corpora encoded within the model parameters devlin2018bert; radford2019language; petroni2019language; (2) contextual knowledge, which is additional information passed into the input prompt kwiatkowski2019natural; joshi2017triviaqa. For certain downstream tasks, such as question-answering, it is essential to augment prompts containing a question/instruction with relevant context for LMs. However, a recent concern is that both the parametric and contextual knowledge may contain private information. Prior work has shown that privacy leakage of parametric knowledge often occurs from memorized pre-training data carlini2019secret. On the other hand, we focus on the privacy leakage of augmented contexts, which can occur when an LM regurgitates them wang2023decodingtrust; priyanshu2023chatbots.
@@
Consider the example shown in Figure 1. An augmented context contains John Doe’s address, and a user queries an LM asking for John Doe’s address. If the output of the LM contains John Doe’s address, then the straightforward approach of comparing the output against the augmented context would suggest there was privacy leakage from the context. This privacy evaluation was performed by prior works that studied data extraction attacks in RAG systems by prompting LMs to regurgitate the context zeng2024good; qi2024follow. However, suppose we re-query the LM but remove (or mask out) John Doe’s address from the context. If the LM still outputs John Doe’s address, then surely the privacy leakage must derive from the LM’s parametric knowledge. Hence, assuming that augmented contexts are not contained in the LM’s parametric knowledge, which may not hold in practice golchin2023time; deng2023investigating; jiang2024investigating, can overestimate the privacy leakage. Indeed, our results in Section 4.2 demonstrates this. Thus, accurately measuring context privacy leakage requires consideration of the existing parametric knowledge of the LM.
Extensive research has examined how factors such as model size, prompt length, and training order contribute to the memorization and subsequent privacy leakage of an LM’s parametric knowledge carlini2021extracting; carlini2022quantifying; biderman2024emergent; lesci2024causal. However, there is a lack of understanding regarding the factors that cause privacy leakage of contextual knowledge. This is challenging as it involves separating the contributions of an LLM’s parametric knowledge from the augmented context longpre2021entity; du2024context, which has implications for solutions that adopt publicly pre-trained LMs to preserve privacy of contexts utpala2023locally; meisenbacher2024dp; flemings2024knowledge.
@@
  • We propose context influence, a metric to principally quantify the privacy leakage of contextual information by measuring the output difference with and without a subset of the context, exemplified in Table 1 with uni-grams. Context influence follows the analysis of differential privacy dwork2006differential, a widely-adopted privacy notion.
  • Then, using a slight reformulation of Context-aware Decoding shi2023trusting, we show that context privacy leakage can be affected (1) explicitly when amplifying/deamplifying contextual knowledge during decoding, and (2) implicitly when contextual knowledge is out-of-distribution with respect to parametric knowledge.
@@ 2.1 Privacy
To understand the privacy leakage of contextual information during decoding, we draw inspiration from Differential Privacy (DP) dwork2006differential; dwork2014algorithmic, a strong privacy notion that gives a provable guarantee on the information leakage. We state the definition below.
Definition 2.1 (Pure Differential Privacy (DP) feldman2021individual).
@@
Definition 2.2 (Ex-post per-instance differential Privacy redberg2021privately).
@@ 2.2 Context-aware Decoding
Satisfying the DP definition requires controlling the amount of privacy that can be leaked by the context during generation. To this end, we borrow ideas from prior work that focused on amplifying contextual information by utilizing Pointwise Mutual Information (PMI) to measure the LM’s dependence on the context, then applying this measurement to the decoding process to explicitly steer the LM’s focus on the context van2022mutual; shi2023trusting. The goal of these prior works does not involve privacy; rather, they are focused on reducing hallucinations by forcing the model to focus more on the context. However, these prior works provide the appropriate foundation to build on for estimating privacy from the context. PMI is defined as
@@
This formulation, known as Context-aware Decoding (CAD) shi2023trusting, helps the LM focus on the context.
@@ 3.1 Motivation
As we argued in the Introduction section, only comparing the LM’s output to the augmented context is insufficient for measuring context privacy leakage. Alternatively, one could estimate the privacy risk by performing Membership Inference Attacks (MIAs) shokri2017membership; jagielski2020auditing. However, this requires instantiating an attacker, which could severely underestimate the privacy leakage, and MIAs have been shown to be oftentimes ineffective on LMs duan2024membership.
Instead, we take a different approach by utilizing the privacy analysis from DP, which provides a strong guarantee that bounds the privacy leakage to any dataset $D$, individual $d_{i}$, and output events $E$. In particular, we follow the observation that sampling-based decoding naturally satisfies the randomized output requirement of differential privacy with respect to the context flemings2024differentially. Moreover, the neighboring definition of DP allows us to separate the contribution from parametric and contextual knowledge, by comparing the output probability distributions with and without subsets of the context. However, the $\epsilon$ bound from Eq. 1 does not provide much insights into the privacy risks of the augmented context, since $\epsilon$ is independent of the context. By using ex-post per-instance DP, we can directly calculate a privacy loss $\epsilon(o,D,D\setminus\{d_{i}\})$ that depends on important parameters– such as the context $D$, neighboring contexts $D\setminus\{d_{i}\}$, generated token oo– which acts as a privacy auditing tool to analyze how these parameters affect the context privacy leakage.
@@ 3.2 Context Influence
[…]
The use of $i$-th token $n$-grams generalizes the privacy level [^1] We restrict the possible substrings to $n$-grams because they still produce natural text (contiguous substrings) while providing a granular measurement of the context influence., where $n=1$ roughly corresponds to word-level privacy xu2020differentially; feyisetan2020privacy while $n=|D|$ corresponds to document-level privacy mattern2022limits; utpala2023locally denoted as $D_{i,n}=D$. Rather than bounding the absolute value of the logs-odds ratio, as is done with DP, definition 3.1 directly uses this quantity as a lower-bound estimate of the privacy leakage of the $i$-th token $n$-gram $D_{i,n}$ for a fixed context $D$ when releasing the next token $y_{t}$. From an adversarial perspective, Definition 3.1 describes how confidently an attacker could infer whether the $i$-th token $n$-gram $D_{i,n}$ is part of the context $D$.
@@
To measure the total context influence over an entire generation $\mathbf{y}$, we simply sum the context influence for each generated token $y_{t}$ which is analogous to the basic composition property of DP dwork2014algorithmic.
@@
One last technicality: context influence only works for sampling-based algorithms, such as temperature sampling ackley1985learning, and not for greedy decoding algorithms, such as argmax. However, top-p holtzman2019curious or top-k sampling fan2018hierarchical can cause potential errors in the context influence calculation unless the selected indices are equal for both $D$ and $D\setminus D_{i,n}$. For the remainder of our work, we focus only on temperature sampling. Specifically, we generate the responses $\mathbf{y}$ to be used for measuring context influence by using a slight reformulation of CAD (Eq. 3), giving us granular control over how much of the contextual knowledge is used during decoding. We introduce Context Influence Decoding (CID) below:
@@
where $\sigma$ is the softmax function and $T$ is the temperature parameter where $T>1$ resulting in a more uniform distribution (i.e. higher entropy) and $0<T<1$ forcing a sharper output distribution.
@@ 3.3 Privacy Leakage with Context Influence
for a fixed $n$. The proof can be found in Appendix B. Meaning regardless of the context $D$, the previously generated tokens $\mathbf{y}_{<t}$, and the next token $y_{t}$, the $i$-th token $n$-gram $D_{i,n}$ influences the LM $\theta$ by at most $\epsilon$; hence, DP bounds the privacy leakage with a context-independent value. For our work, we want to explicitly measure how the privacy leakage changes with respect to the aforementioned variables, hence why chose the analysis of ex-post per-instance DP. This gives a guarantee that the privacy leakage when releasing $y_{t}$ is at least $\tau_{i,n}(\overline{p}_{\theta,\lambda},D,\mathbf{x},\mathbf{y}_{<t},y_{t})$, which follows the more practical direction of privacy auditing jagielski2020auditing.
@@
  • The context $D$ is out-of-distribution with respect to the LM’s parametric knowledge $\theta$, and the subset of contextual information we are interested in is sufficiently large (i.e. large $n$), both maximizing the difference of PMIs. Various factors– including the type of contextual $(D,\mathbf{x})$ and parametric knowledge $\theta$, model size $|\theta|$, and context size |D||D|– can cause the contextual and parametric knowledge to diverge, which we experimentally analyze in Section 4.3. Additionally, in Section 4.4, we compare how different subsets of contextual knowledge influence an LM.
@@
Model Decoding $\lambda$ $\hat{\tau}_{|D|}(\overline{p}_{\theta,\lambda})$
Repeat
Prompts
Rouge
Prompts
$\hat{\tau}_{|D|}(\overline{p}_{\theta,\lambda})$
Repeat
Prompts
Rouge
Prompts
@@ 4.1 Experimental Setup
Datasets. We perform our experimental evaluations on two open-ended generation tasks: summarization via CNN-DM see2017get, a collection of English news articles written by journalists at CNN and the Daily Mail, and long-form question-answering via PubMedQA jin2019pubmedqa, a dataset from the biomedical domain and contexts available. Appendix C contains example prompts used for context influence. Each context document is truncated by $2048$ and $1024$ for PubMedQA and CNN-DM, respectively.
Metrics. We evaluate the generation quality along two dimensions: similarity and faithfulness. For similarity, we employed F1 ROUGE-L lin2004rouge and F1 BERTScore zhang2019bertscore to measure lexical and semantic similarity between the response and the reference, respectively. For faithfulness, we used FactKB feng2023factkb to measure the faithfulness of the response to the context. Our calculation of context influence uses the empirical estimator in Eq. 5 with CID using sampled contexts from CNN-DM and PubMedQA.
Models. Since context influence requires access to the entire model output distribution, we use open-source models due to closed-source ones restricting the output logits. We used OPT 1.3B zhang2022opt, GPT-Neo 1.3B black2021gpt, LLaMA 3 8B and LLaMA 3 8B IT (Instruct) dubey2024llama, and Gemma 2 9B (Instruct) team2024gemma. We set temperature parameter $T=0.8$, the response length to at most 50 tokens, and the number of responses for each dataset is $N=1000$.
@@ 4.2 Context Influence on Input Regurgitation
First, we demonstrate how our context influence metrics offer improvements over directly comparing LLM output with augmented contexts. Following the untargeted attack evaluations by zeng2024good we report:
@@
Next, we show that Repeat Prompts and Rouge Prompts can erroneously indicate privacy leakage from augmented PubMedQA abstracts if the LLM’s parametric knowledge already includes PubMed abstracts. For this, we compare OPT 1.3B and GPT-Neo 1.3B, models with the same number of parameters and similarly follow the GPT-3 architecture brown2020language. GPT-Neo was pre-trained on The Pile dataset gao2020pile, which contains PubMed abstracts. In contrast, OPT was trained on a subset of The Pile that excludes PubMed abstracts zhang2022opt. Therefore, we expect OPT 1.3B to show greater context privacy leakage, as it likely lacks PubMed abstracts in its parametric knowledge and must rely more on the augmented PubMedQA contexts.
@@ 4.3 Factors Contributing to Context Influence
Parametric knowledge $\theta$. We compare context influence on multiple different parametric knowledge sources. Our results indicate that the choice of parametric knowledge can have a substantial affect on the context influence. In particular, we saw from Section 4.2 how just the inclusion of PubMed data in the pre-training data can effectively decrease the context influence of PubMedQA abstracts, creating a false sense of security that GPT-Neo preserves the privacy of the PubMedQA abstracts better than OPT. Hence, this raises an important caveat that smaller context influence does not necessarily imply better context privacy, as one must consider the public data used for pre-training tramer2022considerations.
@@
Context size $|D|$. Additionally, we measured the effect of the context size $|D|$ on context influence for CID using OPT-1.3B. In this setup, we restrict the model to only the first $|D|$ tokens of context for generation and calculating context influence. Shown in Figure 2(b), we observe that when the context is extremely small ($\leq 32)$, then the LM is substantially less influenced by the context. The context may not contain enough relevant information to help the model, and hence, it must rely on its parametric knowledge. However, as we increase the context size from 32 to 256, the model becomes more influenced by the context. After $|D|\geq 256$, the model maintains a relatively constant level of context influence. Hence, truncating the context has marginal affect on context privacy unless the size of the context is substantially reduced.